A consumer with income M, derives utility from consuming good X and Y.
He has a Cobb-Douglas utility function, U(X,Y) = X^1/2 Y^1/2
The prices of good X and Y are Px and Py respectively.
a) Derive his demand functions for X and Y.
b) Derive the price elasticity of demand for X.
c) Is X a normal good?

Need help in part b
1
Expert's answer
2020-11-30T16:41:41-0500
"Solution"

"U(X,Y)=X^{0.5}Y^{0.5}\\"

Solve the consumers optimization problem by maximizing subject to the budget constraint.

"P_x\cdot X+P_y\cdot\ Y\leq\ m\\"

Then, lets use the langrange Theorem to rewrite the constrained optimization problem into a non constrained form,

"\ Max\ L(X, Y, \lambda)={X^{0.5}}{Y^{0.5}}+ \lambda(m-P_xX-P_yY"

The first order condition(necessary) will result in

"0.5XY^{0.5}=\lambda\ P_x......(i)\\
0.5YX^{0.5}=\lambda P_y.......(ii)"

Combining 1 and 2 will result in

"0.5P_yY=0.5P_xX\\
x=\frac{0.5m}{0.5+0.5P_x}\implies is\ the\ demand\ function"

Price elasticity

"=\frac{0.5}{1p}/\frac{0.5}{1p}=1"

c)X is a normal good since it has a positive income elasticity of demand.